# Generalizing subdiffusive Black-Scholes model by variable exponent: Model transformation and numerical approximation

### v1

@@ 5 Numerical experiments
In this section, we provide some numerical examples to verify our theoretical results. We define the following $L^{2}(0,T;L^{2})$ error to measure the accuracy of the numerical solution
$Error_{\tau}(N,M)=\sqrt{\tau\sum_{n=1}^{N}\|U^{2n}-U^{n}\|_{L^{2}}^{2}},\quad\|U^{n}\|_{L^{2}}=\sqrt{h\sum\limits_{j=1}^{M-1}(U^{n}_{j})^{2}},$
$Error_{h}(N,M)=\sqrt{\tau\sum_{n=1}^{N}h\sum_{j=1}^{M-1}\left|U_{2j}^{n}-U_{j}^{n}\right|^{2}},$
@@
$Order_{\tau}=\log_{2}\left(\frac{Error_{\tau}(N,M)}{Error_{\tau}(2N,M)}\right),\quad Order_{h}=\log_{2}\left(\frac{Error_{h}(N,M)}{Error_{h}(N,2M)}\right).$
Example 1. This example considers the case when the exact solution is unknown with $(\bar{a},\bar{b})=(0,1)$, $T=1$, $\sigma=0.45$, $r=0.03$. Set the initial data $u_{0}(x)=\sin(\pi x)$, the forcing term $f=(\frac{r}{\sigma^{4}}-\frac{1}{2\sigma^{2}})\pi\cos(\pi x)$, and $\alpha(t)=\alpha_{0}-\frac{1}{11}t$ $(\alpha^{\prime}(0)\neq 0)$, where $\alpha_{0}\in[0.1,1)$.
Table 5.1: $L^{2}(0,T;L^{2})$-norm errors and temporal convergence orders when $M=32$.
@@
$N$ $Error_{\tau}(N,M)$ $Order_{\tau}$ $N$ $Error_{\tau}(N,M)$ $Order_{\tau}$
16 $1.8337\times 10^{-3}$ 16 $2.0227\times 10^{-3}$
@@
$N$ $Error_{\tau}(N,M)$ $Order_{\tau}$ $N$ $Error_{\tau}(N,M)$ $Order_{\tau}$
16 $4.7357\times 10^{-4}$ 16 $1.3701\times 10^{-4}$
@@
Table 5.2: $L^{2}(0,T;L^{2})$-norm errors and spatial convergence orders when $N=32$.
@@
$M$ $Error_{h}(N,M)$ $Order_{h}$ $M$ $Error_{h}(N,M)$ $Order_{h}$
32 $1.0586\times 10^{-4}$ 32 $9.2023\times 10^{-5}$
@@
$M$ $Error_{h}(N,M)$ $Order_{h}$ $M$ $Error_{h}(N,M)$ $Order_{h}$
32 $7.7674\times 10^{-5}$ 32 $6.5094\times 10^{-5}$
@@
First, with different $\alpha_{0}$, the $L^{2}(0,T;L^{2})$ errors and temporal convergence orders are shown in Table 5.1 when $M=32$. The numerical results in Table 5.1 demonstrate that our method achieves $\frac{1}{2}+\frac{3}{2}\alpha_{0}$ order convergence in the time direction as $N$ increases, which is consistent with the theoretical analysis. In Table 5.2, we explore the $L^{2}(0,T;L^{2})$ error and the order of spatial convergence of the finite element scheme when $N=32$ for different $\alpha_{0}$. As $M$ increases, the second-order convergence rate in space is observed from Table 5.2.
Example 2. This example considers the case when the exact solution is unknown with $(\bar{a},\bar{b})=(0,1)$ and $T=1$. Let the initial data $u_{0}(x)=\sin(\pi x)$, the forcing term $f=0$, $\sigma=0.5$, $r=0.25$ and $\alpha(t)=\alpha_{0}-\frac{1}{11}t^{2}$ ($\alpha^{\prime}(0)=0$), where $\alpha_{0}\in[0.1,1)$.
Here we consider the second example, in which case $\alpha^{\prime}(0)=0$. Table 5.3 reflects the rate of convergence $\approx\frac{1}{2}+\frac{3}{2}\alpha_{0}$ for the finite element scheme at $\alpha_{0}=0.1,0.4,0.7,0.9$, with the order of convergence reaching second order as $\alpha_{0}$ tends to $1$. Table 5.4 lists the $L^{2}(0,T;L^{2})$ errors and the order of spatial convergence for $M=32$. The numerical results show that the finite element method achieves stable second order spatial convergence for all tested values of $\alpha_{0}$, which is consistent with analytical predictions.
Table 5.3: $L^{2}(0,T;L^{2})$-norm errors and temporal convergence orders when $M=32$.
@@
$N$ $E(N,M)$ $Rate^{t}$ $N$ $E(N,M)$ $Rate^{t}$
32 $2.1244\times 10^{-3}$ 32 $1.9582\times 10^{-4}$
@@
$N$ $E(N,M)$ $Rate^{t}$ $N$ $E(N,M)$ $Rate^{t}$
32 $3.6093\times 10^{-4}$ 32 $1.0412\times 10^{-4}$
@@
Table 5.4: $L^{2}(0,T;L^{2})$-norm errors and spatial convergence orders when $N=32$.
@@
$M$ $F(N,M)$ $Rate^{h}$ $M$ $F(N,M)$ $Rate^{h}$
32 $2.4148\times 10^{-4}$ 32 $2.3292\times 10^{-4}$
@@
$M$ $F(N,M)$ $Rate^{h}$ $M$ $F(N,M)$ $Rate^{h}$
32 $2.2778\times 10^{-4}$ 32 $2.2503\times 10^{-4}$
@@
6 Conclusions
In this work, we investigate the subdiffusive Black-Scholes model of variable exponent, which is widely applied in situations where the historical information significantly influences the current pricing. A series of transformations are applied to convert the original model into a feasible formulation such that the numerical methods could be designed and analyzed. Numerical experiments are conducted to confirm the theoretical results. The developed model transformation could also be employed in other option pricing models that will be investigated in the near future.
Acknowledgments
This work was partially supported by the National Social Science Foundation of China under Grant 24BTJ006.

### v2

@@ 5 Numerical experiments
In this section, we provide some numerical examples to verify our theoretical results. A uniform mesh partitioned by the nodal points $x_{i}=ih$ where $0\leq i\leq M$ and $h=\frac{\bar{a}-\bar{b}}{M}$ is used in examples. As the exact solutions are in general not available, we measure the $L^{2}$ error following the two-mesh principle [12, Page 107]
$\text{Error}_{\tau}(N,M)=\sqrt{\tau\sum_{n=1}^{N}h\sum\limits_{j=1}^{M-1}\left|U_{j}^{2n}(2N,M)-U_{j}^{n}(N,M)\right|^{2}},$
$\text{Error}_{h}(N,M)=\sqrt{\tau\sum_{n=1}^{N}h\sum_{j=1}^{M-1}\left|U_{2j}^{n}(N,2M)-U_{j}^{n}(N,M)\right|^{2}},$
@@
$\text{Order}_{\tau}=\log_{2}\left(\frac{\text{Error}_{\tau}(N,M)}{\text{Error}_{\tau}(2N,M)}\right),\quad\text{Order}_{h}=\log_{2}\left(\frac{\text{Error}_{h}(N,M)}{\text{Error}_{h}(N,2M)}\right).$
Example 1. Let $(\bar{a},\bar{b})=(0,1)$, $T=1$, $\sigma=0.45$, $r=0.03$, $u_{0}(x)=\sin(\pi x)$, the forcing term $f=(\frac{r}{\sigma^{4}}-\frac{1}{2\sigma^{2}})\pi\cos(\pi x)$, and $\alpha(t)=\alpha_{0}-\frac{1}{11}t$ $(\alpha^{\prime}(0)\neq 0)$, where $\alpha_{0}\in[0.1,1)$. First, with different $\alpha_{0}$, the $L^{2}$ errors and temporal convergence orders are shown in Table 1 when $M=32$, which demonstrate that the method achieves $\frac{1}{2}+\frac{3}{2}\alpha_{0}$ order convergence in time, which is consistent with the theoretical analysis. In Table 2, we explore the $L^{2}$ error and the order of spatial convergence of the finite element scheme when $N=32$ for different $\alpha_{0}$. As $M$ increases, the second-order convergence rate in space is observed from Table 2.
Table 1: $L^{2}$-norm errors and temporal convergence orders when $M=32$.
@@
NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$ NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$
16 $1.8337\times 10^{-3}$ * 16 $2.0227\times 10^{-3}$ *
@@
NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$ NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$
16 $4.7357\times 10^{-4}$ * 16 $1.3701\times 10^{-4}$ *
@@
Table 2: $L^{2}$-norm errors and spatial convergence orders when $N=32$.
@@
$M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$ $M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$
32 $1.0586\times 10^{-4}$ * 32 $9.2023\times 10^{-5}$ *
@@
$M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$ $M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$
32 $7.7674\times 10^{-5}$ * 32 $6.5094\times 10^{-5}$ *
@@
Example 2. Let $(\bar{a},\bar{b})=(0,1)$, $T=1$, $u_{0}(x)=\sin(\pi x)$, the forcing term $f=0$, $\sigma=0.5$, $r=0.25$ and a nonlinear exponent $\alpha(t)=\alpha_{0}-\frac{1}{11}t^{2}$ ($\alpha^{\prime}(0)=0$), where $\alpha_{0}\in[0.1,1)$. We get the same conclusions as Example 1 from Tables 34.
Table 3: $L^{2}$-norm errors and temporal convergence orders when $M=32$.
@@
NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$ NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$
32 $2.1244\times 10^{-3}$ * 32 $1.9582\times 10^{-4}$ *
@@
NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$ NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$
32 $3.6093\times 10^{-4}$ * 32 $1.0412\times 10^{-4}$ *
@@
Table 4: $L^{2}$-norm errors and spatial convergence orders when $N=32$.
@@
$M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$ $M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$
32 $2.4148\times 10^{-4}$ * 32 $2.3292\times 10^{-4}$ *
@@
$M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$ $M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$
32 $2.2778\times 10^{-4}$ * 32 $2.2503\times 10^{-4}$ *
@@
Example 3. Let $(\bar{a},\bar{b})=(0,1)$, $T=1$, $\sigma=0.4$, $r=0.1$, $u_{0}(x)=\sin(\pi x)$, the forcing term $f=(\frac{r}{\sigma^{4}}-\frac{1}{2\sigma^{2}})\pi\cos(\pi x)$, and $\alpha(t)=\alpha_{0}+\frac{1}{11}t^{3}$ $(\alpha^{\prime}(0)=0)$, where $\alpha_{0}\in(0,0.9]$. Different from the first two examples, the $\alpha(t)$ is an increasing function. We get the same conclusions as Example 1 from Tables 56, which indicates that the monotonicity of $\alpha(t)$ does not affect the numerical accuracy.
Table 5: $L^{2}$-norm errors and temporal convergence orders when $M=32$.
$\alpha_{0}=0.1$ $\alpha_{0}=0.4$
NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$ NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$
4 $3.6526\times 10^{-3}$ * 4 $7.3880\times 10^{-3}$ *
8 $2.3790\times 10^{-3}$ 0.62 8 $3.4375\times 10^{-3}$ 1.10
16 $1.5270\times 10^{-3}$ 0.64 16 $5660\times 10^{-3} 1.1$3
32 $9.7177\times 10^{-4}$ 0.65 32 $7.0100\times 10^{-4}$ 1.16
Theory 0.65 1.10
$\alpha_{0}=0.7$ $\alpha_{0}=0.9$
NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$ NN $\text{Error}_{\tau}(N,M)$ $\text{Order}_{\tau}$
4 $3.3174\times 10^{-3}$ * 4 $1.3359\times 10^{-3}$ *
8 $1.0746\times 10^{-3}$ 1.63 8 $3.5407\times 10^{-4}$ 1.92
16 $3.4713\times 10^{-4}$ 1.63 16 $9.5064\times 10^{-5}$ 1.90
32 $1.1170\times 10^{-4}$ 1.64 32 $2.5463\times 10^{-5}$ 1.90
Theory 1.55 1.85
Table 6: $L^{2}$-norm errors and spatial convergence orders when $N=32$.
$\alpha_{0}=0.1$ $\alpha_{0}=0.4$
$M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$ $M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$
4 $5.3523\times 10^{-3}$ * 4 $4.6527\times 10^{-3}$ *
8 $1.3414\times 10^{-3}$ 2.00 8 $1.1635\times 10^{-3}$ 2.00
16 $3.3542\times 10^{-4}$ 2.00 16 $2.9076\times 10^{-4}$ 2.00
32 $8.3856\times 10^{-5}$ 2.00 32 $7.2682\times 10^{-5}$ 2.00
Theory 2.00 2.00
$\alpha_{0}=0.7$ $\alpha_{0}=0.9$
$M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$ $M$ $\text{Error}_{h}(N,M)$ $\text{Order}_{h}$
4 $3.8724\times 10^{-3}$ * 4 $3.1934\times 10^{-3}$ *
8 $9.6488\times 10^{-4}$ 2.00 8 $7.9319\times 10^{-4}$ 2.01
16 $2.4091\times 10^{-4}$ 2.00 16 $1.9788\times 10^{-4}$ 2.00
32 $6.0207\times 10^{-5}$ 2.00 32 $4.9443\times 10^{-5}$ 2.00
Theory 2.00 2.00
6 Concluding remarks
In this work, we investigate the subdiffusive Black-Scholes model of variable exponent, which is widely applied in situations where the historical information significantly influences the current pricing. A series of transformations are applied to convert the original model into a feasible formulation such that the numerical methods could be designed and analyzed. Numerical experiments are conducted to confirm the theoretical results.
It is worth mentioning that, despite successful applications of the subdiffusive Black-Scholes model of variable exponent [18], the derivation of the variable-exponent model from certain stochastic dynamics requires further investigation. To be specific, it is shown in, e.g. [16], that the subdiffusive Black-Scholes equation with a constant exponent $2H$ for some $0<H<1/2$ could be derived from the stochastic dynamics driven by the fractional Brownian motion $B^{H}_{t}$ with the Hurst index $H$, which is a centered Gaussian process determined by the following covariance function
$\mathbb{E}[B^{H}_{t}\cdot B^{H}_{t}]=\frac{1}{2}\{|t|^{2H}+|s|^{2H}-|t-s|^{2H}\},~~\forall s,t\geq 0.$ (6.1)
Nevertheless, it is pointed out in [3] that “the assumption of a constant Holder regularity (Hurst exponent) in financial time-series seems to be too rigid to address some particularities of a market beyond tranquil period”, and there exist several works considering the time-varying Hurst index in stock markets [2, 6]. In particular, [3] adopts the multifractional Brownian motion $W_{h(t)}$ for some function $h(t)\in(0,1)$, which is also a centered Gaussian process formally defined by its covariance
$\mathbb{E}[W_{h(t)}\cdot W_{h(s)}]=D(h(t),h(s))\{|t|^{h(t)+h(s)}+|s|^{h(t)+h(s)}-|t-s|^{h(t)+h(s)}\},~~\forall s,t\geq 0,$ (6.2)
where
$D(x,y):=\frac{\sqrt{\Gamma(2x+1)\Gamma(2y+1)\sin(\pi x)\sin(\pi y)}}{2\Gamma(x+y+1)\sin(\pi(x+y)/2)},$
to derive a multifractional Black-Scholes model and substantiate its effectiveness in comparison with its counterparts based on standard and fractional Brownian motions. As the exponent in the subdiffusive Black-Scholes model is closely related to the Hurst index as shown above, we adopt the variable exponent in (1.3) following the spirit of the multifractional Brownian motion, and we will work on a rigorous proof from the stochastic dynamics with time-varying Hurst index to the variable-exponent fractional model in the future.
Acknowledgements.
This work was partially supported by the National Social Science Foundation of China under Grant 24BTJ006.Conflict of Interest The authors declare no conflict of interest.